# Iterative Reasoning Preference Optimization

### v2

@@ Abstract
Iterative preference optimization methods have recently been shown to perform well for general instruction tuning tasks, but typically make little improvement on reasoning tasks (Yuan et al., 2024; Chen et al., 2024). In this work we develop an iterative approach that optimizes the preference between competing generated Chain-of-Thought (CoT) candidates by optimizing for winning vs. losing reasoning steps that lead to the correct answer. We train using a modified DPO loss (Rafailov et al., 2023) with an additional negative log-likelihood term, which we find to be crucial. We show reasoning improves across repeated iterations of this scheme. While only relying on examples in the training set, our approach results in increasing accuracy on GSM8K, MATH, and ARC-Challenge for Llama-2-70B-Chat, outperforming other Llama-2-based models not relying on additionally sourced datasets. For example, we see a large improvement from 55.6% to 81.6% on GSM8K and an accuracy of 88.7% with majority voting out of 32 samples.
@@ 1 Introduction
Preference optimization has proven to give large gains when aligning pre-trained language models to human requirements compared to supervised fine-tuning alone (Ziegler et al., 2019; Stiennon et al., 2020). Offline methods such as DPO (Rafailov et al., 2023) are becoming more popular for their simplicity and efficiency. Recent results have shown that iterative application of such an offline procedure is beneficial, whereby the updated model is used to construct new preference relations that are more informative, and hence improve results further. These methods include Iterative DPO (Xu et al., 2023; Xiong et al., 2023), Self-Rewarding LLMs (Yuan et al., 2024), SPIN (Chen et al., 2024), and other methods (Rosset et al., 2024). Common to these approaches is that they have been shown to perform well on general instruction tuning tasks, but they either make only moderate gains or even decrease the performance on standard reasoning tasks. While other kinds of iterative training methods have been applied successfully to reasoning, particularly involving iteration of supervised fine-tuning (SFT) such as STaR (Zelikman et al., 2022), RestEM (Singh et al., 2024), and V-STaR (Hosseini et al., 2024) [^1] V-STaR does use preference optimization, but for training a separate verifier model., using preference optimization to train the generative reasoning model is not applied in these methods.
@@
In this work we develop an approach to apply iterative preference optimization to reasoning tasks, with a particular focus on Chain-of-Thought (CoT) reasoning (Wu et al., 2023). On each iteration we sample multiple chain-of-thought reasoning steps and final answers over training prompts, and then construct preference pairs such that pair winners have correct answers and pair losers have wrong answers. We then train a variant of DPO that includes a negative log-likelihood (NLL) loss term for the pair winners, which also proves crucial for performance. Given the newly trained model, we then iterate the procedure by generating new pairs, and training again, starting from the previously trained model. We find that reasoning performance improves over multiple iterations until it eventually saturates.
@@
We show that our approach, termed Iterative Reasoning Preference Optimization (Iterative RPO), outperforms a number of baselines, including SFT or applying standard DPO, as well as other baselines from the literature. We see an improvement from 55.6% of zero-shot performance on GSM8K to 81.6% after our Iterative RPO training (or from 70.7% to 88.7% with majority voting out of 32 samples), from 77.8% to 86.7% on ARC-Challenge (without using the provided ARC Corpus), and from 12.5% to 20.8% on MATH (without using the provided pretraining corpus in MATH). We provide ablations that indicate the components that lead to these improvements. Overall, our method provides a simple recipe that has the potential to improve the reasoning ability of LLMs over a wide range of tasks.
@@ 2 Iterative Reasoning Preference Optimization
Initialization
We assume we are given an initial model $M_{0}$, and a training set $D=\{x_{i},y_{i}\}$ containing questions $x_{i}$ and their correct answers $y_{i}$ The model will be trained and updated at each iteration, resulting in models $M_{0},M_{1},\dots,M_{T}$.
Chain-of-Thought & Answer Generation
@@
$(c_{i}^{n},y_{i}^{n})\sim M_{t}(x_{i})\quad\text{for all}\ x_{i}\in D\ \text{and}\ n\in[1,N].$
@@
$G_{i}^{w}=\{c_{i}^{n},y_{i}^{n}\,|\,r_{i}^{n}=1\}\quad\text{and}\quad G_{i}^{l}=\{c_{i}^{n},y_{i}^{n}\,|\,r_{i}^{n}=0\}.$
Next we build a dataset of preference pairs by selecting a winner response $(c_{i}^{w},y_{i}^{w})$ from $G_{i}^{w}$, and a loser response $(c_{i}^{l},y_{i}^{l})$ from $G_{i}^{l}$. In particular, we simply iterate over $G_{i}^{w}$ and $G_{i}^{l}$ simultaneously [^2] If the iteration reaches the end of a set, it restarts from the first element. If one of the sets is empty, then that input will be ignored. to produce $K$ pairs $\{w_{k},l_{k}\}$, in order to ensure we use as much of the data as possible.
$D_{t}^{\text{pairs}}=\{(c_{i}^{w_{k}},y_{i}^{w_{k}}),(c_{i}^{l_{k}},y_{i}^{l_{k}})\ |\ \text{for all}\ x_{i}\in D\text{ and }k\in[1,K]\}.$
@@
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{NLL}}(c_{i}^{w},y_{i}^{w}|x_{i})+\alpha\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})$
$=-\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{|c_{i}^{w}|+|y_{i}^{w}|}-\alpha\log\sigma\left(\beta\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{\log M_{t}(c_{i}^{w},y_{i}^{w}|x_{i})}-\beta\frac{\log M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{\log M_{t}(c_{i}^{l},y_{i}^{l}|x_{i})}\right).$ (1)
Here $M(x)$ denotes the probability of sequence $x$ under the model $M$, and $\sigma$ is the sigmoid function. We use the previous iteration’s model $M_{t}$ as the reference model in the denominator of the DPO term. Note that the NLL term is normalized by the total response length. The hyperparameter $\alpha$ balances the two loss terms. For brevity we omit the pair index $k$, but we optimize this loss on each of the $k\in[1,K]$ pairs generated for every input sample. At the end of this training, we thus obtain our next model $M_{t+1}=M_{\theta}$, which will be then used to build data for the subsequent iteration.
Iterative Training
@@ 3.1 Math Word Problems: GSM8K
In our first set of experiments, we use the GSM8K dataset (Cobbe et al., 2021) that contains real grade-school math word problems. Each problem contains a question $x_{i}$, gold chain-of-thought solution $c_{i}$, and a final numerical answer $y_{i}$. For our entire training process, we only use the training set of around 7.5k problems without any extra data.
As a seed model $M_{0}$ we use the chat version of Llama-2 70B model (Touvron et al., 2023), which is instruction fine-tuned. We use a zero-shot prompt containing the question together with instructions to produce a chain-of-thought and to follow a specific format so the final answer can be easily extracted (the exact prompt is given in subsection A.1). In each iteration, we generate $N=30$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.3 for iterations 3–4 (hoping that there is a significant number of incorrect generations in later iterations). Since some problems might not have any model-generated correct solution, we include the gold human written solution $(c_{i},y_{i})$ in the winning set $G_{i}^{w}$ so it is not empty. Then we generate $K=10$ pairs per problem for training with our loss in Equation 1, and filtered out examples that were too long in terms of overflowing the context length or else do not have any incorrect generations. This gave around 55–60k pairs for training, per iteration. [^3] If after filtering the number of pairs is larger than 60k, then we randomly select around 60k examples. This number is fixed because we do not want to introduce another source of variability in our experiments.
In total, we performed four iterations, producing models $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. For each iteration, we train a maximum of 5000 steps, then select the best checkpoint using a held-out 1k samples from the training set. We then retrain including those 1k samples for the selected number of steps. The coefficient $\alpha$ is tuned in {0.5, 1, 2, 4} when training M1, and we end up using 1 for all experiments in the paper. The coefficient $\beta$ in the DPO loss is tuned in {0.05, 0.1, 0.5, 1.0}, and we end up using 0.1 in this experiment. We used a batch size of 16 and a learning rate 7e-7.
@@
Iterative RPO improves over baselines
We find that Iterative RPO outperforms zero-shot CoT, supervised fine-tuning (SFT) on the gold CoT solutions, and variants of DPO by a wide margin. SFT gives a boost in performance compared to zero-shot CoT from 55.6% to 63.5% but still far from the 81.6% of Iterative RPO. We apply standard DPO to the same set of preference pairs $D_{0}^{\text{pairs}}$ as used in the first iteration of our method. Whether initializing from Llama-2-70b-chat ($M_{0}$) or from SFT training on the chosen (winner) examples, we find that DPO performance, while being better than zero-shot CoT, is no better than the SFT model, with accuracies of 61.8% or 60.3% respectively. We also show that SFT on only the chosen CoT solutions, which corresponds to the first iteration of the STaR method, improves results to 65.2% over SFT on the gold solutions alone, but still falls short of the performance of the first iteration of Iterative RPO. One hypothesis for these improvements is the necessity of including the rejected sequences in the training objective, otherwise their probability increases along with the chosen samples, see Figure 2. We note this observation has also been reported in concurrent work (Hong et al., 2024). All of the results reported above are using a single generation at test time using greedy decoding. If we use majority voting over 32 samples (sampling with temperature 0.8), a standard approach to improve performance in the literature, we can improve the accuracy of our approach from 81.1% to 88.2% for iteration 3, and from 81.6% to 88.7% for iteration 4 of Iterative RPO.
Iterations of Iterative RPO yield improved reasoning
We observe that Iterative RPO provides improvements over its training iterations, increasing the base model accuracy by 47% (from 55.6% to 81.6%) in total. In contrast, supervised training using the gold CoT only brings about a 14% accuracy boost. We see performance improves across each iteration, from 73.1% to 78.0% to 81.1% to 81.6%. However, the gain decays across the iterations (17.5%, 4.9%, 3.1%, 0.5%), indicating an upper limit on learning across iterations, especially as we are iterating across a fixed number of prompts, i.e., only from the training samples. We also show that it is the iterations of updating the model (i.e., initializing from the previous model) that are helping, not just because there is more data in the form of new pairs generated from the fixed training set. To test this statement, we run the first iteration of Iterative RPO but on twice as much paired data by doubling $K$, as well as the STaR method first iteration with twice as much data as well. In both cases performance improves compared to less data, but not as much as performing two iterations. Iterative RPO with twice as much data obtains 74.8% (an improvement over 73.1% using the original dataset size); however, training for two iterations obtains 78.0%. For STaR, training on twice as much data obtains 66.9%, compared to 65.2% with the original data, which is still a much lower performance than Iterative RPO.
@@
Figure 2: Effect of SFT training. (a) Although SFT training (solid green) is on chosen sequences ($D_{0}^{\text{pairs}}$, from iterative RPO iteration 1) only, the rejected sequence log probabilities also increase and are close to the chosen sequence probabilities. In contrast, our DPO+NLL training (blue) manages to decrease the rejected probabilities while increasing the chosen probabilities. This observation could potentially help explain why SFT-only performance lags significantly behind Iterative RPO Iteration 1 performance. (b) We show a similar plot but where SFT is trained on gold (dataset-provided) CoTs. Chosen and rejected sequence probabilities (which are from $D_{0}^{\text{pairs}}$) are still close to each other, but with a slightly bigger gap. In addition, the chosen sequence probabilities barely increase.
@@
Figure 3: Effect of NLL loss term on DPO training. In our GSM8K experiments we observe the log probability of chosen sequences in standard DPO without NLL loss (solid orange) decreases over training steps, especially if the model is initialized from SFT training on chosen sequences (right). However, they increase over training steps when using DPO with NLL loss (solid blue). In all four settings, the margin between the two curves continues increasing. We find that DPO+NLL loss gives superior test accuracy in our experiments.
NLL loss is necessary in our method: DPO with NLL vs. DPO without NLL
The first iteration of our method can be compared to standard DPO training, which uses the same preference data, as reported in Table 1. We see a large performance drop (73.1% vs. 61.8%) using DPO compared to our method after 1 iteration. The gap remains large even when the standard DPO training starts from the superior SFT-tuned model, which it has been argued improves DPO’s performance (Rafailov et al., 2023, 2024). Our results support the need of the NLL loss term in our training, not just using SFT for initialization. To further understand this, we plot the sequence-level log probability over training steps for these methods in Figure 3. We see that for DPO without NLL loss there is a decrease over training for the chosen sequences, whereas for DPO with NLL there is not, which may help explain the improved performance of the latter. We note that related observations have been made elsewhere in various settings (Pal et al., 2024; Xu et al., 2024; Hong et al., 2024). Further, we note that whether we initialize with Llama-2-70b-chat or SFT on chosen for Iterative RPO, accuracy results of first iteration training do not seem to deviate (both obtain the same score 73.1%).
Other results in the literature
We can compare our results to others in the literature, even if their experiments are in different settings. Touvron et al. (2023) reports an accuracy of 56.8% for 8-shot Llama-2-70b, which is close to our zero-shot CoT results for Llama-2-70b-chat. In terms of closed-source proprietary language models, some results are superior results to ours, while others are not, for example GPT-4 obtains 92.0% (5-shot chain-of-thought) (Achiam et al., 2023), Claude 2 obtains 88.0% (Anthropic Team, 2023), PaLM 2 obtains 80.7% (Anil et al., 2023), while GPT-3.5 obtains 57.1% (5-shot) (Achiam et al., 2023). We note that the size (number of parameters) and the makeup of the training set of some of these models have not been fully disclosed. For results that use the same size and class model, Llama-2-70b, MetaMath (Yu et al., 2023) reports an accuracy of 82.3%, while WizardMath reports 81.6% (Luo et al., 2023). These last two results use additional augmented training data, whereas our method does not use additional prompts. We note that such approaches should be orthogonal to ours, and both can provide benefits.

### v3

@@ Abstract
Iterative preference optimization methods have recently been shown to perform well for general instruction tuning tasks, but typically make little improvement on reasoning tasks (Yuan et al., 2024; Chen et al., 2024). In this work we develop an iterative approach that optimizes the preference between competing generated Chain-of-Thought (CoT) candidates by optimizing for winning vs. losing reasoning steps. We train using a modified DPO loss (Rafailov et al., 2023) with an additional negative log-likelihood term, which we find to be crucial. We show reasoning improves across repeated iterations of this scheme. While only relying on examples in the training set, our approach results in increasing accuracy on GSM8K, MATH, and ARC-Challenge for Llama-2-70B-Chat, outperforming other Llama-2-based models not relying on additionally sourced datasets. For example, we see a large improvement from 55.6% to 81.6% on GSM8K and an accuracy of 88.7% with majority voting out of 32 samples.
@@ 1 Introduction
Preference optimization has proven to give large gains when aligning pre-trained language models to human requirements compared to supervised fine-tuning alone (Ziegler et al., 2019; Stiennon et al., 2020). Offline methods such as DPO (Rafailov et al., 2023) are becoming more popular for their simplicity and efficiency. Recent results have shown that iterative application of such an offline procedure is beneficial, whereby the updated model is used to construct new preference relations that are more informative, and hence improve results further. These methods include Iterative DPO (Xu et al., 2023; Xiong et al., 2023), Self-Rewarding LLMs (Yuan et al., 2024), SPIN (Chen et al., 2024), and other methods (Rosset et al., 2024). Common to these approaches is that they have been shown to perform well on general instruction tuning tasks, but they either make only moderate gains or even decrease the performance on standard reasoning tasks. While other kinds of iterative training methods have been applied successfully to reasoning, particularly involving the iteration of supervised fine-tuning (SFT) such as STaR (Zelikman et al., 2022), RestEM (Singh et al., 2024), and V-STaR (Hosseini et al., 2024) [^1] V-STaR does use preference optimization, but for training a separate verifier model., using preference optimization to train the generative reasoning model is not applied in these methods.
@@
In this work, we develop an approach to apply iterative preference optimization to reasoning tasks, with a particular focus on Chain-of-Thought (CoT) reasoning (Wu et al., 2023). On each iteration we sample multiple chain-of-thought reasoning steps and final answers over training prompts, and then construct preference pairs such that pair winners have correct answers and pair losers have wrong answers. We then train a variant of DPO that includes a negative log-likelihood (NLL) loss term for the pair winners, which also proves crucial for performance. Given the newly trained model, we then iterate the procedure by generating new pairs, and training again, starting from the previously trained model. We find that reasoning performance improves over multiple iterations until it eventually saturates.
@@
We show that our approach, termed Iterative Reasoning Preference Optimization (Iterative RPO), outperforms a number of baselines, including SFT or applying standard DPO, as well as other baselines from the literature. We see an improvement from 55.6% of zero-shot performance on GSM8K to 81.6% after our Iterative RPO training (or from 70.7% to 88.7% with majority voting out of 32 samples), from 77.8% to 86.7% on ARC-Challenge (without using the provided ARC Corpus), and from 12.5% to 20.8% on MATH (or from 18.8% to 29.1% with majority voting out of 32 samples), without using the provided pretraining corpus in MATH. We provide ablations that indicate the components that lead to these improvements. We also present analysis on how different objectives influence the probabilities of training sequences, which helps explain the success of our method. Overall, our method provides a simple recipe that has the potential to improve the reasoning ability of LLMs over a wide range of tasks, as shown on the three tasks we consider.
@@ 2 Iterative Reasoning Preference Optimization
Initialization.
We assume we are given an initial model $M_{0}$, and a training set $D=\{(x_{i},y_{i})\}_{i}$ containing questions $x_{i}$ and their correct answers $y_{i}$. The model will be trained and updated at each iteration, resulting in models $M_{0},M_{1},\dots,M_{T}$.
Chain-of-thought & answer generation.
@@
$(c_{i}^{n},y_{i}^{n})\sim M_{t}(x_{i})\quad\text{for all}\ x_{i}\in D\ \text{and}\ n\in[N],$
where we use $[N]$ to denote $\{1,2,\dots,N\}$.
@@
$G_{i}^{w}=\{c_{i}^{n},y_{i}^{n}\,|\,r_{i}^{n}=1\},$
$G_{i}^{l}=\{c_{i}^{n},y_{i}^{n}\,|\,r_{i}^{n}=0\}.$
Next we build a dataset of preference pairs by selecting a winner response $(c_{i}^{w},y_{i}^{w})$ from $G_{i}^{w}$, and a loser response $(c_{i}^{l},y_{i}^{l})$ from $G_{i}^{l}$. In particular, we simply iterate over $G_{i}^{w}$ and $G_{i}^{l}$ simultaneously [^2] If the iteration reaches the end of a set, it restarts from the first element. If one of the sets is empty, then that input will be ignored. to produce $K$ pairs of indices $\{(w_{k},l_{k})\}$, in order to ensure we use as much of the data as possible.
$D_{t}^{\text{pairs}}=\{(c_{i}^{w_{k}},y_{i}^{w_{k}}),(c_{i}^{l_{k}},y_{i}^{l_{k}})\ |\ \text{for all}\ x_{i}\in D\text{ and }k\in[K]\}.$
@@
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})+\alpha\mathcal{L}_{\text{NLL}}(c_{i}^{w},y_{i}^{w}|x_{i})$
$=-\log\sigma\left(\beta\log\frac{M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{M_{t}(c_{i}^{w},y_{i}^{w}|x_{i})}-\beta\log\frac{M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{M_{t}(c_{i}^{l},y_{i}^{l}|x_{i})}\right)-\alpha\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{|c_{i}^{w}|+|y_{i}^{w}|}.$ (1)
Here $M(x)$ denotes the probability of sequence $x$ under the model $M$, and $\sigma$ is the sigmoid function. We use the previous iteration’s model $M_{t}$ as the reference model in the denominator of the DPO term. Note that the NLL term is normalized by the total response length. The hyperparameter $\alpha$ balances the two loss terms. For brevity we omit the pair index $k$, but we optimize this loss on each of the $k\in[K]$ pairs generated for every input sample. At the end of this training, we thus obtain our next model $M_{t+1}=M_{\theta}$, which will be then used to build data for the subsequent iteration.
Iterative training.
@@
  • $M_{4}$
    : Initialized with $M_{3}$, then trained with $D_{3}^{\text{pairs}}$ using $\mathcal{L}_{\text{DPO+NLL}}$.
@@ 3.1 Math Word Problems: GSM8K
In our first set of experiments, we use the GSM8K dataset (Cobbe et al., 2021) [^3] We have confirmed that the licenses of the datasets used in this paper (MIT for GSM8K and MATH, CC BY-SA 4.0 for ARC) are respected. that contains real grade-school math word problems. For example the question: “Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?”. These questions typically require the model to perform intermediate reasoning, i.e. generating chain-of-thought before answering, otherwise performance is poor. Each problem contains a question $x_{i}$, gold chain-of-thought solution $c_{i}$, and a final numerical answer $y_{i}$. For our entire training process, we only use the training set of around 7.5k problems without any extra questions.
Experimental setup.
As a seed model $M_{0}$ we use the chat version of Llama-2 70B model (Touvron et al., 2023), which is instruction fine-tuned. We use a zero-shot prompt containing the question together with instructions to produce a chain-of-thought and to follow a specific format so the final answer can be easily extracted (the exact prompt is given in subsection B.1). In each iteration, we generate $N=30$ solutions per problem using sampling with temperature 0.8 for iterations 1–2 and temperature 1.3 for iterations 3–4 (hoping that there is a significant number of incorrect generations in later iterations). Since some problems might not have any model-generated correct solution, we include the gold human written solution $(c_{i},y_{i})$ in the winning set $G_{i}^{w}$ so it is not empty. Then we generate $K=10$ pairs per problem for training with our loss in Equation 1, and filter out examples that were too long in terms of overflowing the context length or else do not have any incorrect generations. This procedure gives around 55–60k pairs for training, per iteration. [^4] If after filtering the number of pairs is larger than 60k, then we randomly select around 60k examples. This number is fixed because we do not want to introduce another source of variability in our experiments.
In total, we perform four iterations, producing models $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. For each iteration, we train a maximum of 5000 steps, and then select the best checkpoint using a held-out 1k samples from the training set. We then retrain while including those 1k samples for the selected number of steps. The coefficient $\alpha$ is tuned in {0.25, 0.5, 1, 2} when training $M_{1}$, and we end up using 1 for all experiments in the paper. The coefficient $\beta$ in the DPO loss is tuned in {0.05, 0.1, 0.5, 1.0}, and we end up using 0.1 in this experiment. We use a batch size of 16 and a learning rate 7e-7 using the AdamW optimizer. Throughout this paper, all generation is done using one node containing eight V100 GPUs (32G memory). All training is done using eight nodes each containing eight A100 GPUs (80G memory).
@@
Figure 2: Effect of SFT training. (a) Although SFT training (solid green) is on chosen sequences ($D_{0}^{\text{pairs}}$, from iterative RPO iteration 1) only, the rejected sequence log probabilities (dotted green) also increase and are close to the chosen sequence probabilities. In contrast, our DPO+NLL training (blue) manages to decrease the rejected probabilities while increasing the chosen probabilities. This observation could potentially help explain why SFT-only performance lags significantly behind Iterative RPO Iteration 1 performance. (b) We show a similar plot but where SFT is trained on gold (dataset-provided) CoTs. Chosen and rejected sequence probabilities (which are from $D_{0}^{\text{pairs}}$) are still close to each other, but with a slightly bigger gap. Another observation is that the chosen sequence probabilities barely increase.
Iterative RPO improves over baselines.
We find that Iterative RPO outperforms zero-shot CoT, supervised fine-tuning (SFT) on the gold (dataset-provided) CoT solutions, and variants of DPO by a wide margin. SFT gives a boost in performance compared to zero-shot CoT from 55.6% to 63.5% but still far from the 81.6% of Iterative RPO. We apply standard DPO to the same set of preference pairs $D_{0}^{\text{pairs}}$ as used in the first iteration of our method. Whether initializing from Llama-2-70b-chat ($M_{0}$) or from SFT training on the chosen (winner) examples, we find that DPO performance, while being better than zero-shot CoT, is no better than the SFT model, with accuracies of 61.8% or 60.3% respectively.
We also show that SFT on only the chosen CoT solutions, which corresponds to the first iteration of the STaR method, improves results to 65.2% over SFT on the gold solutions alone, but still falls short of the performance of the first iteration of Iterative RPO. One hypothesis for these improvements is the necessity of including the rejected sequences in the training objective; otherwise their probability increases along with the chosen samples; see Figure 2. We note this observation has also been reported in concurrent work (Hong et al., 2024).
All of the results reported above are using a single generation at test time using greedy decoding. If we use majority voting over 32 samples (sampling with temperature 0.8), a standard approach to improve performance in the literature, we can improve the accuracy of our approach from 81.1% to 88.2% for iteration 3, and from 81.6% to 88.7% for iteration 4 of Iterative RPO. While performance is much improved using majority vote, this should be compared to a majority vote baseline, where we find a similarly large improvement over the zero-shot chain-of-thought with majority vote, which obtains an accuracy of 70.7%.
Iterations of Iterative RPO yield improved reasoning.
We observe that Iterative RPO provides improvements over its training iterations, increasing the base model accuracy by 47% (from 55.6% to 81.6%) in total. In contrast, supervised training using the gold CoT only brings about a 14% accuracy boost. We see performance improves across each iteration, from 73.1% to 78.0% to 81.1% to 81.6%. However, the gain decays across the iterations (17.5%, 4.9%, 3.1%, 0.5%), indicating an upper limit on learning across iterations, especially as we are iterating across a fixed number of prompts, i.e., only from the training samples.
We also show that it is the iterations of updating the model (i.e., initializing from the previous model) that are helping, not just because there is more data in the form of new pairs generated from the fixed training set. To test this statement, we run the first iteration of Iterative RPO but on twice as much paired data by doubling $K$, and we run the STaR method first iteration with twice as much data as well. In both cases performance improves compared to less data, but not as much as performing two iterations. Iterative RPO with twice as much data obtains 74.8% (an improvement over 73.1% using the original dataset size); however, training for two iterations obtains 78.0%. For STaR, training on twice as much data obtains 66.9%, compared to 65.2% with the original data, which is still a much lower performance than Iterative RPO.
@@
Figure 3: Effect of NLL loss term on DPO training for GSM8K. In our GSM8K experiments we observe the log probability of chosen sequences in standard DPO without NLL loss (solid orange) decreases over training steps, especially if the model is initialized from SFT training on chosen sequences (right). However, they increase over training steps when using DPO with NLL loss (solid blue). In all four settings, the margin between the two curves continues increasing. We find that DPO+NLL loss gives superior test accuracy in our experiments.
NLL loss is necessary in our method: DPO with NLL vs. DPO without NLL.
The first iteration of our method can be compared to standard DPO training, which uses the same preference data, as reported in Table 1. We see a large performance drop (73.1% vs. 61.8%) using DPO compared to our method after one iteration. The gap remains large even when the standard DPO training starts from the superior SFT-tuned model, which it has been argued improves DPO’s performance (Rafailov et al., 2023, 2024). Our results support the need of the NLL loss term in our training, not just using SFT for initialization. To further understand this, we plot the sequence-level log probability over training steps for these methods in Figure 3. We see that for DPO without NLL loss there is a decrease over training for the chosen sequences, whereas for DPO with NLL there is not, which may help explain the improved performance of the latter. We note that related observations have been made elsewhere in various settings (Pal et al., 2024; Xu et al., 2024; Hong et al., 2024). Further, we note that whether we initialize with Llama-2-70b-chat or SFT on chosen for Iterative RPO, accuracy results of first iteration training do not seem to deviate (both obtain the same score 73.1%). This is another advantage of our method as the training process is simpler without the SFT step.
Other results in the literature.
We can compare our results to others in the literature, even if their experiments are in different settings. Touvron et al. (2023) reports an accuracy of 56.8% for 8-shot Llama-2-70b, which is close to our zero-shot CoT results for Llama-2-70b-chat. In terms of closed-source proprietary language models, some results are superior to ours, while others are not; for example GPT-4 obtains 92.0% (5-shot chain-of-thought) (Achiam et al., 2023), Claude 2 obtains 88.0% (Anthropic Team, 2023), PaLM 2 obtains 80.7% (Anil et al., 2023), while GPT-3.5 obtains 57.1% (5-shot) (Achiam et al., 2023). We note that the size (number of parameters) and the makeup of the training set of some of these models have not been fully disclosed. For results that use the same size and class model, Llama-2-70b, MetaMath (Yu et al., 2024) reports an accuracy of 82.3%, while WizardMath reports 81.6% (Luo et al., 2023). These last two results use additional augmented training data, whereas our method does not use additional prompts. Such approaches should be orthogonal to ours, and both can provide benefits.